OpenStudy (anonymous):

What is the equation of the quadratic graph with a focus of (8, −8) and a directrix of y = −6?

OpenStudy (anonymous):

@Venny can you help me?

OpenStudy (anonymous):

Since the focus is at (8,-8) we know that the center will have an x coordinate of 8. Now, the center will be halfway between the focus and directrix which will have a y coordinate of 7, so we know the center is at (8,-7) Now we can see that the center is 1 unit away from both the focus and directrix, so p=1. This also opens downward because the focus is lower than the directrix. \((x - h)^2 =-4p(y - k)\) Here h=8, k=-7 and p=1 \((x-8)^2=-4(y+7)\)

OpenStudy (anonymous):

ok i did all that and my final answer was -1/4(x-8)^2-7

OpenStudy (anonymous):

That's right, just dont forget the y=, :)

OpenStudy (anonymous):

ok thank you so much! (: